# Query complexities of quantum channel discrimination and estimation: A unified approach

### v1

@@ 8.2 Lower bounds on query complexities of channel estimation
In this section, we leverage the lower bounds from (3.30)–(3.31), along with our lower bounds on the error probabilities and query complexities of channel discrimination from Section 8.1, in order to establish lower bounds on the minimax error probabilities and query complexities of channel estimation. We then establish asymptotic lower bounds on these quantities in terms of SLD Fisher information, arriving at results that generalize those from [MKSF+25, Eq. (53)].
@@
For a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)$ of channels, the following bounds hold for the minimax error probability of channel estimation:
$1-4p_{e,\|}\left(1-p_{e,\|}\right)$ $\leq n\inf_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\\ \theta,\theta^{\prime}\in\Theta,\\ W\in\mathbb{L}:\left\|W\right\|\leq 1\end{subarray}}\left\{a_{W}^{\theta,\theta^{\prime}}+\left(n-1\right)b_{W}^{\theta,\theta^{\prime}}\right\},$ (8.44)
$1-4p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)$ $\leq n\inf_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\\ \theta,\theta^{\prime}\in\Theta,\\ W\in\mathbb{L}:\left\|W\right\|\leq 1\end{subarray}}\left\{a_{W}^{\theta,\theta^{\prime}}+\left(n-1\right)\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}\right\},$ (8.45)
@@
For a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)$ of channels and for all $\varepsilon\in\left(0,\frac{1}{2}\right]$, the following lower bounds hold for the query complexities of channel estimation:
$n_{\|}^{*}(\varepsilon,\delta,\mathcal{F})$ $\geq\left\lceil\sup_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta,\\ W\in\mathbb{L}:\left\|W\right\|\leq 1\end{subarray}}\frac{b_{W}^{\theta,\theta^{\prime}}-a_{W}^{\theta,\theta^{\prime}}+\sqrt{\left(b_{W}^{\theta,\theta^{\prime}}-a_{W}^{\theta,\theta^{\prime}}\right)^{2}+4b_{W}^{\theta,\theta^{\prime}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2b_{W}^{\theta,\theta^{\prime}}}\right\rceil$ (8.51)
$\geq\left\lceil\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{W\in\mathbb{L}}\left\{a_{W}^{\theta,\theta^{\prime}}:\left\|W\right\|\leq 1,\ M_{W}^{\theta,\theta^{\prime}}=I,\ \left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta\right\}}\right\rceil,$ (8.52)
$n_{\operatorname{ad}}^{*}(\varepsilon,\delta,\mathcal{F})$ $\geq\left\lceil\sup_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta,\\ W\in\mathbb{L}:\left\|W\right\|\leq 1\end{subarray}}\frac{c_{W}^{\theta,\theta^{\prime}}+\sqrt{\left(c_{W}^{\theta,\theta^{\prime}}\right)^{2}+4\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}}\right\rceil$ (8.53)
$\geq\left\lceil\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{W\in\mathbb{L}}\left\{a_{W}^{\theta,\theta^{\prime}}:\left\|W\right\|\leq 1,\ M_{W}^{\theta,\theta^{\prime}}=I,\ \left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta\right\}}\right\rceil,$ (8.54)
@@
Similar to what was stated in Remark 26, the lower bounds in (8.52) and (8.54) are only applicable in the case that, for some $\theta,\theta^{\prime}\in\Theta$ satisfying $\left|\theta-\theta^{\prime}\right|>2\delta$, there exists $W\in\mathbb{L}$ such that $M_{W}^{\theta,\theta^{\prime}}=I$. Otherwise these lower bounds are trivially equal to 0.
Finally, we obtain asymptotic expressions for bounds on the minimax error probabilities and query complexities by appealing to (8.12), (8.14), (3.30), (3.31), and Theorems 15, 21, and 22.
@@
$n_{\|}^{*}(\varepsilon,\delta,\mathcal{F})$ $\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{b_{\theta}\right\}}}\right)$ (8.61)
@@
$n_{\operatorname{ad}}^{*}(\varepsilon,\delta,\mathcal{F})$ $\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\sqrt{a_{\theta}b_{\theta}}\right\}}}\right)$ (8.63)
@@
$n_{\|}^{*}(\varepsilon,\delta,\mathcal{F})\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{b_{\theta}\right\}}}\right),$ (8.74)
@@
$n_{\operatorname{ad}}^{*}(\varepsilon,\delta,\mathcal{F})\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\sqrt{a_{\theta}b_{\theta}}\right\}}}\right),$ (8.81)
@@
Similar to the approaches outlined in Section 8.1.1, lower bounds on the query complexities of channel estimation can be efficiently computed by a combination of semi-definite optimization and grid search. Indeed, from the bounds in Corollary 30 and algorithms similar to Algorithms 28 and 29, combined with an outer search over $\theta,\theta^{\prime}\in\Theta$ satisfying $\left|\theta-\theta^{\prime}\right|>2\delta$, one can efficiently compute lower bounds on the query complexities of channel estimation. One can obtain looser bounds at lower computational cost by performing an outer search over just $\theta\in\Theta$ and setting $\theta^{\prime}=\theta+2\delta$. For sufficiently small $\delta$, one could alternatively obtain lower bounds on the query complexities of channel estimation using the bounds from Corollary 33 and algorithms similar to Algorithms 28 and 29, combined with an outer search over $\theta\in\Theta$.
@@ 9.2 Optimization of quantities based on Bures distance of channels
$\sqrt{F}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\inf_{\rho}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]$ (9.18)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\inf_{\rho}\operatorname{Tr}\!\left[\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\rho\right]$ (9.19)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sup_{\lambda\in\mathbb{R}}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I\right\}$ (9.20)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sup_{\lambda\geq 0}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I\right\}$ (9.21)
@@
$d_{B}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})=\inf_{\begin{subarray}{c}\lambda\in\mathbb{R},\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda:\lambda I\geq 2\left(I-\Re\!\left[M\right]\right)\geq-\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\\ M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\end{array}\right\}.$ (9.23)
@@
$d_{B}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\left\|2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\right\|$ (9.24)
@@
$\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\left\{\left\|2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\right\|+\left(n-1\right)\left\|I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right\|^{2}\right\}\\ =\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda+\left(n-1\right)\mu:\\ \lambda I\geq 2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\geq-\lambda I,\\ \begin{bmatrix}\mu I&I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\\ I-\left(V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right)^{{\dagger}}&I\end{bmatrix}\geq 0,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0\end{array}\right\},$ (9.29)
@@ 10 Conclusion
Acknowledgements—ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and an RMIT Vice Chancellor’s Senior Research Fellowship. JJM acknolwedges support from QuantERA (HQCC). She is also grateful to the Cornell School of Electrical and Computer Engineering for hospitality during an October 2025 research visit, as well as the Cornell Lab of Ornithology for the Merlin app, which has enabled many hours of fruitful birding. TN acknowledges the support from the Department of Mathematics and the IQUIST Postdoctoral Fellowship from the Illinois Quantum Information Science and Technology Center at the University of Illinois Urbana-Champaign. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.

### v2

@@ 8.2 Lower bounds on query complexities of channel estimation
In this section, we leverage the lower bounds from (3.30)–(3.31), along with our lower bounds on the error probabilities and query complexities of channel discrimination from Section 8.1, in order to establish lower bounds on the minimax error probabilities and query complexities of channel estimation. We then establish asymptotic lower bounds on these quantities in terms of SLD Fisher information, arriving at results that generalize those from [MKSF+25, Eq. (53)] (see (8.61)–(8.64)). These asymptotic lower bounds bear similarity to quantum Cramer–Rao bounds for channel estimation, which have appeared in the context of quantum estimation theory (see, e.g., [DDKG12, KDD13, DDM14, ZJ21, KGADD23]).
@@
For a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of channels, the following bounds hold for the minimax error probability of channel estimation:
$1-4p_{e,\|}\left(1-p_{e,\|}\right)$ $\leq n\inf_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B}\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\left\{a_{W}^{\theta,\theta^{\prime}}+\left(n-1\right)b_{W}^{\theta,\theta^{\prime}}\right\},$ (8.44)
$1-4p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)$ $\leq n\inf_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B}\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\left\{a_{W}^{\theta,\theta^{\prime}}+\left(n-1\right)\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}\right\},$ (8.45)
@@
For a parameterized family $\mathcal{F}\equiv\left(\mathcal{N}_{\theta}\right)_{\theta\in\Theta}$ of channels and for all $\varepsilon\in\left(0,\frac{1}{2}\right]$, the following lower bounds hold for the query complexities of channel estimation:
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\left\lceil\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B},\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\frac{b_{W}^{\theta,\theta^{\prime}}-a_{W}^{\theta,\theta^{\prime}}+\sqrt{\left(b_{W}^{\theta,\theta^{\prime}}-a_{W}^{\theta,\theta^{\prime}}\right)^{2}+4b_{W}^{\theta,\theta^{\prime}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2b_{W}^{\theta,\theta^{\prime}}}\right\rceil$ (8.51)
$\geq\left\lceil\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{W\in\mathbb{B}}\left\{a_{W}^{\theta,\theta^{\prime}}:M_{W}^{\theta,\theta^{\prime}}=I,\ \left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta\right\}}\right\rceil,$ (8.52)
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\left\lceil\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta,W\in\mathbb{B},\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}\frac{c_{W}^{\theta,\theta^{\prime}}+\sqrt{\left(c_{W}^{\theta,\theta^{\prime}}\right)^{2}+4\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}\left(1-4\varepsilon\left(1-\varepsilon\right)\right)}}{2\sqrt{a_{W}^{\theta,\theta^{\prime}}b_{W}^{\theta,\theta^{\prime}}}}\right\rceil$ (8.53)
$\geq\left\lceil\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{W\in\mathbb{B}}\left\{a_{W}^{\theta,\theta^{\prime}}:M_{W}^{\theta,\theta^{\prime}}=I,\ \left|\theta-\theta^{\prime}\right|>2\delta,\ \theta,\theta^{\prime}\in\Theta\right\}}\right\rceil,$ (8.54)
@@
Similar to what was stated in Remark 27, the lower bounds in (8.52) and (8.54) are only applicable in the case that, for some $\theta,\theta^{\prime}\in\Theta$ satisfying $\left|\theta-\theta^{\prime}\right|>2\delta$, there exists $W\in\mathbb{B}$ such that $M_{W}^{\theta,\theta^{\prime}}=I$. Otherwise these lower bounds are trivially equal to 0.
Finally, we obtain asymptotic expressions for bounds on the minimax error probabilities and query complexities by appealing to (8.12), (8.14), (3.30), (3.31), and Theorems 16, 22, and 23.
@@
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{b_{\theta}\right\}}}\right)$ (8.61)
@@
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\sqrt{a_{\theta}b_{\theta}}\right\}}}\right)$ (8.63)
@@
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{b_{\theta}\right\}}}\right),$ (8.74)
@@
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})\geq O\!\left(\frac{1}{\delta}\sqrt{\frac{1-4\varepsilon\left(1-\varepsilon\right)}{\inf_{\theta\in\Theta,\left(H_{\theta}\right)_{\theta\in\Theta}}\left\{\sqrt{a_{\theta}b_{\theta}}\right\}}}\right),$ (8.81)
@@
Similar to the approaches outlined in Section 8.1.1, lower bounds on the query complexities of channel estimation can be efficiently computed by a combination of semi-definite optimization and grid search. Indeed, from the bounds in Corollary 31 and algorithms similar to Algorithms 29 and 30, combined with an outer search over $\theta,\theta^{\prime}\in\Theta$ satisfying $\left|\theta-\theta^{\prime}\right|>2\delta$, one can efficiently compute lower bounds on the query complexities of channel estimation. One can obtain looser bounds at lower computational cost by performing an outer search over just $\theta\in\Theta$ and setting $\theta^{\prime}=\theta+2\delta$. For sufficiently small $\delta$, one could alternatively obtain lower bounds on the query complexities of channel estimation using the bounds from Corollary 34 and algorithms similar to Algorithms 29 and 30, combined with an outer search over $\theta\in\Theta$.
@@ 9.2 Optimization of quantities based on Bures distance of channels
$\sqrt{F}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\sup_{W\in\mathbb{B}}\inf_{\rho}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]$ (9.18)
$=\sup_{W\in\mathbb{B}}\inf_{\rho}\operatorname{Tr}\!\left[\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\rho\right]$ (9.19)
$=\sup_{W\in\mathbb{B}}\sup_{\lambda\in\mathbb{R}}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I\right\}$ (9.20)
$=\sup_{W\in\mathbb{B}}\sup_{\lambda\geq 0}\left\{\lambda:\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\geq\lambda I\right\}$ (9.21)
@@
$d_{B}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})=\inf_{\begin{subarray}{c}\lambda\in\mathbb{R},W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda:\lambda I\geq 2\left(I-\Re\!\left[M\right]\right)\geq-\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\\ M=V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\end{array}\right\}.$ (9.23)
@@
$d_{B}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})$ $=\inf_{W\in\mathbb{B}}\left\|2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\right\|$ (9.24)
@@
$\inf_{W\in\mathbb{B}}\left\{\left\|2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\right\|+\left(n-1\right)\left\|I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right\|^{2}\right\}\\ =\inf_{\begin{subarray}{c}\lambda,\mu\geq 0,\\ W\in\mathbb{L}\end{subarray}}\left\{\begin{array}[]{c}\lambda+\left(n-1\right)\mu:\\ \lambda I\geq 2\left(I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right)\geq-\lambda I,\\ \begin{bmatrix}\mu I&I-V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\\ I-\left(V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right)^{{\dagger}}&I\end{bmatrix}\geq 0,\\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0\end{array}\right\},$ (9.29)
@@ 10 Conclusion
Acknowledgements—ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and an RMIT Vice Chancellor’s Senior Research Fellowship. She is also grateful to the Cornell School of Electrical and Computer Engineering for hospitality during an October 2025 research visit, as well as the Cornell Lab of Ornithology for the Merlin app, which has enabled many hours of fruitful birding. JJM acknowledges support from QuantERA (HQCC). TN acknowledges support from the Department of Mathematics and the IQUIST Postdoctoral Fellowship from the Illinois Quantum Information Science and Technology Center at the University of Illinois Urbana-Champaign. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.